OpenStudy (anonymous):

When integrating with respect to y to find the area of a region, how do you know which function is the upper boundary?

OpenStudy (7thgrader):

Make sure you didn't give any medals on a post yet, and open the same post in a new tab, then give a medal to two different people.

OpenStudy (anonymous):

What???

OpenStudy (kainui):

Let's say you're integrating y=x^2 from 0 to 4 on y, then it could look like this 

OpenStudy (anonymous):

Yes

OpenStudy (kainui):

Really all we're doing is taking y=x^2 and renaming the y and x to be the others like this x=y^2 and solving for our new "y" as y=sqrt(x) and integrating from 0 to 4. See, not really doing anything too crazy I hope.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But how do you determine which one is the upper or lower boundary? I tried flipping the paper and going with whichever one is on top, but that ALWAYS gives me the wrong one

OpenStudy (anonymous):